I need to understand how to do percentage:
1.4% of 28
6 is what percent of 25
12 is 6% if (?).
4 1/4% of (?) = 8.5
also: 2(m+1) + 3 = (m-2)5
I just need to understand these concepts.
Thanks

Answer by tutorcecilia(2152)

For percentages, use the following formula:
Part/Whole = (%)/100 [Plug-in the known values and solve for the unknown values]
"of" =multiplication
Try to identify which is the "whole" and which is the "part"
.
1.4% of 28
Part/Whole = (%)/100
Part/28=1.4/100 [solve for the "Part"]
x/28=1.4/100 [cross-multiply]
100x=(28)(1.4)
100x=39.2
x=39.2/100
x=.392
Checking by plugging all of the values back into the original equation and solve each sid of the equation. both sides shoule match:
.
Checking:
Part/Whole = (%)/100
.392/28=1.4/100
0.014=1.4/100
0.014=0.014 [checks out]
.
6 is what percent of 25
Part/Whole = (%)/100
6/25=x/100 [6 is the "part"; 25 is the "whole"; the "percent" is missing]
(6)(100)=25x [solve for x% by cross-multiplying]
600=25x
600/25=x
24=x or 24%
.
Checking:
Part/Whole = (%)/100
6/25=24/100
0.24=.24 [checks out]
.
12 is 6% if (?).
Part/Whole = (%)/100
12/Whole=6/100 [the "whole" is missing]
12/x=6/100 [solve for x by cross-multiplying]
(12)(100)=(6%(x)
1200=x6
1200/6=x6/6
200=x
.
Checking:
Part/Whole = (%)/100
12/200=6/100
0.06=0.06 [checks out]

4 1/4% of (?) = 8.5
4.25% of x=8.5 [change the fraction to a decimal]
Part/Whole = (%)/100 [the "whole" is missing]
8.5/x=4.25/100
(8.5)(100)=(4.25)(x)
850=4.25x
850/4.25=4.25x/4.25
200=x
.
Checking:
Part/Whole = (%)/100
8.5/200=4.25/100
0.0425=0.0425 [checks out]
.
2(m+1) + 3 = (m-2)5
2(m+1) + 3 = (m-2)5 [use the distributive property to solve]
2m+2 +3=5m-10 [solve for the m=term]
2m+5=5m-10
2m-2m+5=5m-2m-10
5=3m-10
5+10=3m-10+10
15=3m
15/3=3/m
5=m
Check by plugging (m=5) back into the original equation and solve]
2(m+1) + 3 = (m-2)5
2((5)+1) + 3 = ((5)-2)5
2(6)+3=(3)(5)
12+3=15
15=15 [checks out]